# Wilsonian Renormalization of Neural Network Gaussian Processes

### v2

@@ Appendix A Spectral decomposition of the Gaussian action
$S=$ $\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (81)
$+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$
@@
$=\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}$
×$\qquad\times\left(\sum_{k}f_{mk}\phi_{k}(x)\right)K^{-1}(x,x^{\prime})\left(\sum_{k^{\prime}}f_{mk^{\prime}}\phi_{k^{\prime}}(x^{\prime})\right)$
@@
$=\int\!\mathrm{d}\mu_{x}\left[\sum_{k}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]$
×$\qquad\quad\times\left[\sum_{k^{\prime}}\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\phi_{k^{\prime}}(x)\right]$
@@
$\int$ $\mathrm{d}\mu_{x}\left(\sum_{k\leq\kappa}f_{mk}\phi_{k}+\sum_{k>\kappa}f_{mk}\phi_{k}\right)$ (84)
×$\qquad\times\left(\sum_{k^{\prime}\leq\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k^{\prime}}+\sum_{k^{\prime}>\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k}^{\prime}\right)$
$=\left(\sum_{k,k^{\prime}\leq\kappa}+\sum_{k,k^{\prime}>\kappa}+\sum_{\begin{subarray}{c}k\leq\kappa\\ k^{\prime}>\kappa\end{subarray}}+\sum_{\begin{subarray}{c}k>\kappa\\ k^{\prime}\leq\kappa\end{subarray}}\right)\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
@@
Appendix B Simplifying interactions in the replica limit
@@
$C_{mm^{\prime}}\rightarrow\sum_{k>\kappa}\left(f_{km}f_{km^{\prime}}+y_{k}(f_{km}+f_{km^{\prime}})\right)$ (86)
@@
Appendix C $\delta W_{>}(x)$ dependent terms
@@
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (87)
×$\qquad\qquad\times\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
@@
As we are keeping only leading order in $W$, under the GP integration this term would simply be replaced by its average under the free greater mode action where it would yield the following additional contribution to the renormalized GP action
$C_{1}\coloneqq$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$ (88)
×$\qquad\qquad\times\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$
Next we consider the term proportional to $\frac{\delta W^{(1)}_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$. As it is even in the greater feature modes, we have
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (89)
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}$
$\qquad\qquad-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}^{3}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\left(\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)$
$\qquad\qquad-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\left(3\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)$
@@
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (90)
×$\qquad\qquad\times\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$
@@
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$ (91)
×$\qquad\qquad\times\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$
@@
where the first term in the first line vanishes since $He_{2}(\phi_{q}^{2})$ integrates to zero under the standard normal measure. This yeilds the following contribution to the renormalized GP action
$C_{5}\coloneqq$ $-\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$ (92)
×$\qquad\qquad\times\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right]^{2}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$

### v3

@@ Appendix A Spectral decomposition of the Gaussian action
$S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (73)
@@
$=\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\left(\sum_{k}f_{mk}\phi_{k}(x)\right)K^{-1}(x,x^{\prime})\left(\sum_{k^{\prime}}f_{mk^{\prime}}\phi_{k^{\prime}}(x^{\prime})\right)$
@@
$=\int\!\mathrm{d}\mu_{x}\left[\sum_{k}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]\left[\sum_{k^{\prime}}\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\phi_{k^{\prime}}(x)\right]$
@@
$\int$ $\mathrm{d}\mu_{x}\left(\sum_{k\leq\kappa}f_{mk}\phi_{k}+\sum_{k>\kappa}f_{mk}\phi_{k}\right)\left(\sum_{k^{\prime}\leq\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k^{\prime}}+\sum_{k^{\prime}>\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k}^{\prime}\right)$ (76)
$=\left(\sum_{k,k^{\prime}\leq\kappa}+\sum_{k,k^{\prime}>\kappa}+\sum_{\begin{subarray}{c}k\leq\kappa\\ k^{\prime}>\kappa\end{subarray}}+\sum_{\begin{subarray}{c}k>\kappa\\ k^{\prime}\leq\kappa\end{subarray}}\right)\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
@@
Appendix B Applying the Woodbury identity
In this short appendix, we detail the application of the Woodbury matrix identity in (28). In particular, we apply the identity in the form
$\mathbb{1}_{M\times M}-U\left(\mathbb{1}_{N\times N}+VU\right)^{-1}V=\left(\mathbb{1}_{M\times M}+UV\right)^{-1}\leavevmode\nobreak\ ,$ (77)
where $U$ is $M\!\times\!N$ and $V$ is $N\!\times\!M$. Since the matrix ${C=FF^{\top}}$ defined in (27) has dimensions $M\!\times\!M$, we identify $U=F/\sigma$ and $V=F^{\top}/\sigma$, whereupon (77) becomes
$\mathbb{1}-\frac{1}{\sigma^{2}}F\left(\mathbb{1}+F^{\top}F/\sigma^{2}\right)^{-1}F^{\top}=\left(\mathbb{1}+FF^{\top}/\sigma^{2}\right)^{-1}\leavevmode\nobreak\ .$ (78)
A trivial rearrangement yields
$F$ $\left(\mathbb{1}+F^{\top}F/\sigma^{2}\right)^{-1}F^{\top}=\sigma^{2}-\sigma^{2}\left(\mathbb{1}+C/\sigma^{2}\right)^{-1}$ (79)
$=\sigma^{2}-\frac{\sigma^{4}}{\sigma^{2}+C}=\frac{\sigma^{2}C}{\sigma^{2}+C}=\left(C^{-1}+\sigma^{-2}\right)^{-1}\leavevmode\nobreak\ ,$
which is (28).
Appendix C Simplifying interactions in the replica limit
@@
$C_{mm^{\prime}}\rightarrow\sum_{k>\kappa}\left(f_{km}f_{km^{\prime}}-y_{k}(f_{km}+f_{km^{\prime}})\right)$ (81)
@@
Appendix D $\delta W_{>}(x)$ dependent terms
@@
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$ (82)
@@
As we are keeping only the leading order in $W$, under the GP integration this term would simply be replaced by its average under the free greater mode action where it would yield the following additional contribution to the renormalized GP action
$C_{1}\coloneqq\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$ (83)
Next, we consider the term proportional to $\frac{\delta W^{(1)}_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$. As it is even in the greater feature modes, we have
-- $\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$ (84)
$=-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\delta W^{(1)}_{>}\varphi_{q^{\prime}}^{3}$
$=-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\left(\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)-2\left[\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q^{\prime}}\right]^{3}\left(3\sum_{k_{1}\neq q^{\prime}}W_{k_{1},q^{\prime}}\varphi_{k_{1}}\right)$
@@
-$\int$- ∫ $\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\frac{\delta W^{(1)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$ (85)
@@
-$\int$- ∫ $\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)\frac{\delta W^{(2)}_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}$ (86)
@@
where the first term in the first line vanishes since $He_{2}(\phi_{q}^{2})$ integrates to zero under the standard normal measure. This yields the following contribution to the renormalized GP action
$C_{5}\coloneqq-\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right]^{2}\frac{W_{q^{\prime}q^{\prime}}\lambda_{q^{\prime}}}{\sigma^{2}}\leavevmode\nobreak\ .$ (87)
@@
Appendix E Effect of varying threshold $T$ in real-world data experiments
In this appendix we include additional plots in figure 4 that show the effect of the choice of threshold, $T$, on the numerical experiments presented in section IV.1.
Figure 4: Effect of varying the learnability threshold, $T$. For the two real-world dataset examples (MNIST and CIFAR10) considered in section  IV.1, we show the effect of the choice of learnability threshold $T\in(0,1)$. The choice of $T$ inadvertently determines the learnability cutoff, $\kappa$, through finding $\kappa$ such that the learnability factor $L_{\kappa}\approx T$ (see equation 37). While this choice slightly changes the RG theory prediction at high $\eta$, the effect is relatively minor.
Appendix F Exploration of the distribution of feature modes as a function of $k$
In figure 2 we saw that the feature modes for real-world datasets, like the ones considered in section IV.1 (MNIST and CIFAR10), are often Gaussian distributed. This interesting fact has been previously asserted in the literature Simon2021 , motivating Gaussianity as a starting point for theoretical descriptions. In this appendix, we explore the dependence of this statement as a function of $k$ for the real-world datasets considered in this work. Namely, we can see in figure 5 that as we go to very high $k$ values, the distribution of feature modes become better described by a Cauchy distribution.
Figure 6 further explores this trend; for each $k$ we perform both a Gaussian and Cauchy fit to the feature mode distribution and analyze the quality of the fit by calculating the one-sided KS test statistic between the empirical data and best fit. We note that for the majority of the $k$ range, a Gaussian distribution remains a better approximation, as indicated by having a lower KS test statistic value. We additionally calculate the inverse participation ratio, $p_{k}$ THOULESS197493 , which is defined as
$p_{k}:=\sum_{i}\frac{[\varphi_{k}(x_{i})]^{4}}{\left(\sum_{j}[\varphi_{k}(x_{j})]^{2}\right)^{2}}\leavevmode\nobreak\ .$ (88)
We can see that these Cauchy-like modes correspond to more data-point-localized features (as indicated by having a larger $p_{k}$). Moreover, these Cauchy-like modes describe well the bulk of the probability for very high $k$ which corresponds to very small $\lambda_{k}$. Thus, the contribution to $\sigma_{\rm eff}^{2}$ from these very high ($k>8000$) modes is $\mathcal{O}(1\%)$. Overall, we can see that a Gaussian approximation of feature modes is well-motivated but still incomplete; the presence of such non-Gaussian features motivates studying corrections to this approximation, as has been done in this work (see section V.)
Figure 5: Example distribution of high-$k$ feature modes. By comparing with figure 2, we can see that the feature modes shift from being best described by a Gaussian distribution to being better described by a Cauchy distribution for high $k$ values. The best fit Gaussian and best fit Cauchy distributions are overlaid.
Figure 6: Distribution of feature modes as a function of $k$. Here we show that feature modes shift from being best described by a Gaussian distribution to being better described by a Cauchy distribution at high $k$ values. For each mode, we perform a fit to both a Gaussian and Cauchy distribution and plot the one-sided KS test statistic between the empirical data and best fit as a function of $k$. Indeed, for very high $k$, we can see that a Cauchy distribution becomes a better description. By also showing the inverse participation ratio, $p_{k}$ THOULESS197493 , we see that these Cauchy-like modes correspond to more data-point-localized features (as indicated by having a larger inverse participation ratio).